Tesla stock currently sells for $30 a share. The stock will pay a per-share dividend of $3 in one year. The dividend is expected to grow at a constant rate of 8% per year. a. Calculate the stock price one year from now. b. Calculate the dividend yield (DY) for Year 1. c. Calculate the capital gains yield (CGY) for Year 1. d. Calculate the stock price two years from now.
Answer #1

r = D1/Po + g

r = 3/30 + 8% = 18%

a) Dividend at time t = 2 equals 3*(1+8%)

Price = 3*(1+8%)/(.18 - .08) = 32.40

b) Dividend at year 1 = 3*(1+8%)/32.40 = 10%

c) Capital gains yield = growth rate = 8%

d) Stock price two year from now

3*(1+8%)^2/(.18-.08) = 34.99